# Self-Supervised Pre-Training with Equilibrium Constraints

### v1

@@ V-B Multilingual Pre-Training
TABLE IV: Multilingual speech data used in self-supervised pre-training and downstream fine-tuning/test tasks.
We pre-train a Conformer model using raw audio. For feature extraction, we employ a 1D convolutional layer with a kernel size of 3, which is used in both pre-training and fine-tuning to capture local temporal dependencies in the input signal before passing it to the Conformer encoder. During pre-training, the CPC loss is computed using a context length of 256 samples, along with 12 positive and 12 negative samples per instance. The model consists of 8 Conformer blocks, each containing 512 hidden units and 8 attention heads, where each head has a 64-dimensional representation. The convolutional kernel size used within the Conformer blocks is 31. The total number of parameters is 59M.
When training the CPC baseline, we start the training with a learning rate of 5e35e-3 and continue training for 80 epoch which is then annealed by $\frac{1}{10}$ every epoch for next 20 epochs. The training ends at 100 epochs. When training with the proposed PTEC, we start the training with a local learning rate $\alpha\!=\!5e-3$ and global learning rate $\beta\!=\!5e-4$ for 60 epochs. They are then annealed by $\frac{1}{10}$ every epoch afterward synchronously. The training ends after 80 epochs. During PTEC training, the lower level optimization on each language is also conducted in parallel. All the training uses the AdamW optimizer.
After pre-training, a linear classification layer is added to the pre-trained Conformer model. The model is then fine-tuned on labeled speech data for each downstream ASR task using the CTC criterion. The softmax layer comprises 1,000 labels, which are generated using SentencePiece [26]. Fine-tuning is performed for 30 epochs with an initial learning rate of $\beta\!=\!5\!\times\!10^{-5}$, which is reduced by a factor of 10 after each subsequent epoch. The fine-tuning process concludes after 50 epochs.
@@
PTEC.3 37.6 47.8 70.1 55.2
TABLE V: WERs of iterative CSSL and PTEC on four downstream fine-tuning ASR tasks on different languages. In the training, CSSL and PTEC alternatingly initialize each other.
Table V compares the WERs of CSSL, used as the baseline, and the proposed PTEC across four downstream ASR tasks in different languages. Based on insights from multi-domain training, we adopt a one-step local update ($K=1$) for the multilingual experiments. We also follow the same iterative mutual initialization process as in the multi-domain experiments. We conduct three rounds of sequential training, where CSSL is trained for 100 epochs, followed by 80 epochs of PTEC training in each round. The results from the multilingual experiments exhibit a similar trend to those observed in the multi-domain setting. The proposed iterative training strategy significantly enhances the adaptivity of the pre-trained model, leading to consistent WER reductions across all four unseen languages. Specifically, when comparing the final WERs after iterative training (the last row) to the baseline (the first row), we observe relative improvements of 2.9%–9.6% across the four test sets.
VI Summary
In this paper we propose PTEC, a new self-supervised pre-training strategy, to deal with data heterogeneity by imposing equilibrium constraints from various data sources after $K$-step gradient descent. These constraints are used to preserve local characteristics of feature representation in addition to the averaged global loss in the conventional pre-training approaches. We use the first-order approximation to solve the proposed bilevel optimization problem and discuss its connection to MAML. Experiments are carried out on multi-domain and multilingual ASR tasks. It shows that the proposed PTEC strategy can significantly improve the adaptivity of the pre-trained model which can yield improved performance in downstream fine-tuning tasks.

### v2

@@ 5.2 Multilingual Pre-Training
Table 4: Multilingual speech data used in self-supervised pre-training and downstream fine-tuning/test tasks.
We pre-train a Conformer model using raw audio. We employ a 1D convolutional layer with a kernel size of 3 to capture local temporal dependencies in the input signal before passing it to the Conformer encoder. During pre-training, the CPC loss is computed using a context length of 256 samples, along with 12 positive and 12 negative samples per instance. The model consists of 8 Conformer blocks, each containing 512 hidden units and 8 attention heads of 64 dimensions. The convolutional kernel size is 31. The total number of parameters is 59M.
For the CPC baseline, we start the training with a learning rate of 5e35e\!-\!3 and continue training for 80 epochs which is then annealed by $\frac{1}{10}$ every epoch for next 20 epochs. The training ends at 100 epochs. For PTLOC, we start the training with a local learning rate $\alpha\!=\!5e\!-\!3$ and global learning rate $\beta\!=\!5e\!-\!4$ for 60 epochs. They are then annealed by $\frac{1}{10}$ every epoch afterward synchronously. The training ends after 80 epochs. All the training uses the AdamW optimizer.
After pre-training, a linear classification layer is added to the pre-trained Conformer model for FT on labeled data for each downstream ASR task using the CTC criterion. The softmax layer comprises 1,000 labels, which are generated using SentencePiece [25]. FT is performed for 30 epochs with an initial learning rate of $\beta\!=\!\!5e\!-\!5$, which is reduced by a factor of 10 after each subsequent epoch. The FT concludes after 50 epochs.
@@
PTLOC.3 37.6 47.8 70.1 55.2
Table 5: WERs of iterative CSSL and PTLOC on four downstream fine-tuning ASR tasks on different languages. In the training, CSSL and PTLOC alternately initialize each other.
Table 5 compares the WERs of CSSL, used as the baseline, and the proposed PTLOC across four downstream ASR tasks in different languages. We also follow the same iterative mutual initialization process as in the multi-domain experiments. We conduct three rounds of sequential training, where CSSL is trained for 100 epochs, followed by 80 epochs of PTLOC training in each round. The results from the multilingual experiments exhibit a similar trend to those observed in the multi-domain setting. The proposed iterative training strategy significantly enhances the adaptivity of the pre-trained model, leading to consistent WER reductions across all four unseen languages. Specifically, when comparing the final WERs after iterative training (the last row) to the baseline (the first row), we observe relative improvements of 2.9%–9.6% across the four test sets.
6 Summary
In this paper, we propose PTLOC to deal with data heterogeneity in self-supervised pre-training. Local constraints are imposed to ensure that the models optimize each data source to its local optimum after K-step gradient descent initialized from the model. We use the first-order approximation to solve the resulting bilevel optimization problem. Experiments are carried out on multi-domain and multilingual ASR tasks. It shows that PTLOC can significantly improve the adaptivity of the pre-trained model, which can yield improved performance in downstream fine-tuning tasks.